Rate Law for Two Reactant Systems For the reaction the rate law can be written as: Rate = k[A]rm[B]n where k is the rate constant for the reaction. Under the condition [B]。>> [A]o , a good linear regression fit can be obtained for a plot of [(1)/([A 1 (M-1) versus t (s). For two sets of differing initial concentrations, the equations of were found to be: Equation of regression line[Blo y = 2.09% 103x + 7.46× 103 1.25 M y = 4.10x10% + 5.81×10-1 1.75 M What is the value of m? m= What is the value of n? n= 1 pts Submit Answer Tries 0/99 What is the rate constant k (with consistent units)? k=Please help with these 3 questions. I know that it is second order
Answer #1

In regression line equation, 1/[A] is y and t is x. Now since the reaction follows second order rate law,

\frac{1}{[A]}=\frac{1}{[A_{o}]}+kt

Solving the two equations, we get,

0=((4.10-2.09)\times10^{3})x+(5.81-7.46)\times10^{3}

x=\frac{1.65}{2.01}=3=m

y=(2.09\times10^{3}\times3)+(7.46\times10^{3})

y=13730=k

Now,

m+n=2

3+n=2

n=-1